OpenStudy (anonymous):

Factor fully: 2m²+m-15

OpenStudy (anonymous):

[-1(+/-)2 root(30)] / 4

OpenStudy (anonymous):

u can do it with quadratic equations..,

OpenStudy (anonymous):

2m²+m-15 (2m - 5) (m + 3) m1 = 5/2, and m2 = -3

OpenStudy (anonymous):

@gerryliyana how did you come up with these coefficients? like what is the method to solve non-monic quadratics

OpenStudy (anonymous):

I was looking for how much value when multiplied result is 15, and when combined the result 1